# How to Train Neural Field Representations A Comprehensive Study and Benchmark

### v1

@@ 7 Conclusion
The goal of this work is to gain a deeper understanding of NeFs as data representations. To this end, we developed a library for fast generation and manipulation of Neural Datasets for different modalities, such as images and shapes. The speed gain in generating Neural Datasets is substantial compared to naive approaches, practically enabling the use of neural fields for representing datasets at scale.
The study we conducted resulted in a few significant insights into what constitutes better quality NeF representations. The first key takeaway is that obtaining good quality NeF representations is a trade-off between good reconstruction and how spread NeFs are in parameter space. Generally, to achieve good reconstruction more gradient descent iterations are needed, which would spread the NeF representations in parameter space, negatively affecting their quality. Because of that, we propose to fit Neural Datasets from the same shared initialization, and demonstrate that this simple practice results in significantly higher representation quality across different modalities. Second, overfitting hurts representation quality. In other words, good reconstruction helps downstream performances up until the NeFs are overfitting the signals. To address this, we propose to stop fitting the representations early, before overfitting happens.
@@
Limitations We limited our study to images and shapes leaving out neural radiance fields (NeRFs) and videos despite our library being suited to those types of NeFs. Additionally, for practical reasons, we choose to measure representation quality of NeFs using the classification accuracy of a GNN classifier. Although classification is often used in the representation learning literature to evaluate representation quality [4, 10, 27], we acknowledge that good classification accuracy might not result in good downstream performance for general tasks. We plan to explore other downstream tasks and include them in the Neural Field Arena in future work.
8 Acknowledgements
This work was funded by ROV with grant number PPS2102 and Elekta Oncology AB and was supported by an institutional grant of the Dutch Cancer Society and the Dutch Ministry of Health.
We thank SURF (www.surf.nl) for the support in using the National Supercomputer Snellius.
@@ 1 NeFs architectures
Metrics for reconstruction quality
@@ 4 Extent of the study
In the initial phase, we searched a large space of hyperparameters using a smaller subset of the whole dataset, and trained the classifier only for $10$ epochs. In the initial hyperparameter exploration, we looked at optimization – see Table 4 – and architecture hyperparameters – see Table 5. For this exploration, we used the TPE sampler implemented in Optuna [5, 2]. In short, this sampler uses the previous results to inform the choice of the best hyperparameters to use next. Each experiment took between a few minutes to a few hours to run, depending on the number of steps and the size of the architecture. This was done for all $4$ datasets and $3$ models.
In the second phase, we ran a full grid search over the initialization scheme, the number of steps, and the hidden dimension of the NeFs. These were the most impactful parameters according to the search performed in the first phase. This grid search was performed for all datasets and models. Refer to Table 6.

### v2

@@ 7 Conclusion
The goal of this work is to gain a deeper understanding of NeFs as data representations. To this end, we developed a library for fast generation and manipulation of Neural Datasets for different modalities, such as images and shapes. The speed gain in fitting Neural Datasets is substantial compared to naive approaches, practically enabling the use of neural fields for representing datasets at scale.
The study we conducted resulted in significant insights into what constituted better-quality NeF representations. The first key takeaway is that obtaining good NeF representations is a trade-off between good reconstruction and how spread NeFs are in parameter space. Generally, to achieve good reconstruction more gradient descent iterations are needed, which would spread the NeF representations in parameter space, negatively affecting their quality. Because of that, we propose to fit Neural Datasets from the same shared initialization, and demonstrate that this simple practice results in significantly higher representation quality across different modalities. Second, overfitting hurts representation quality. In other words, good reconstruction helps downstream performances up until the NeFs are overfitting the signals. To address this, we propose to stop fitting the representations early, before overfitting happens. Finally, we show how an increase in the number of parameters used to fit a signal leads to a decrease in downstream performance.
@@
Limitations. We limited our study to images and shapes leaving out neural radiance fields (NeRFs) and videos despite our library being suited to those types of NeFs. Additionally, for practical reasons, we choose to measure representation quality of NeFs using the classification accuracy of a GNN classifier. Although classification is often used in the representation learning literature to evaluate representation quality [4, 10, 29], we acknowledge that good classification accuracy might not result in good downstream performance for general tasks. We plan to explore other downstream tasks and include them in the Neural Field Arena in future work.
@@ 1 NeFs Architectures
Metrics for reconstruction quality.
@@ 3 The Classifier Architectures
Figure 9: Training and test accuracy for three different-size GNN and MLP models on a CIFAR10 neural dataset. Overfitting ability is clear also for small capacities.
3.1 Ablation on the Architectures
To evaluate the ability of the models to overfit to the neural representations, we carried out an ablation study using both an MLP and a GNN. The MLP uses linear layers with ReLU activations and batch normalization [22]. The GNN is the one previously described. For both, we decide to train a small/base/large versions. For the MLP these correspond to 74k/116k/363k parameters, while for the GNN they correspond to 500k/1.2M/1.5M parameters. In both cases, we use a CIFAR10 dataset from our study and simply use the neural representations as inputs. For the MLP the input is the concatenated and flattened version of the weights and biases, while for the GNN it is the actual graph of the neural field, where the connectivity is determined by the weights, and node features by the biases. We show the training and validation curves in Fig. 9. These show clear signs of overfitting, the training curve of the GNN slowly approaches the maximum accuracy, while the one of the MLP does so quickly. For both, the validation curves are almost identical at all scales, likely due to capacity saturation.
@@ 4 Extent of the Study
In the initial phase, we searched a large space of hyperparameters using a smaller subset of the whole dataset, and trained the classifier only for $10$ epochs. In the initial hyperparameter exploration, we looked at optimization – see Tab. 4 – and architecture hyperparameters – see Tab. 5. For this exploration, we used the TPE sampler implemented in Optuna [5, 2]. In short, this sampler uses the previous results to inform the choice of the best hyperparameters to use next. Each experiment took between a few minutes to a few hours to run, depending on the number of steps and the size of the architecture. This was done for all $4$ datasets and $3$ models.
In the second phase, we ran a full grid search over the initialization scheme, the number of steps, and the hidden dimension of the NeFs. These were the most impactful parameters according to the search performed in the first phase. This grid search was performed for all datasets and models. Refer to Tab. 6.